# Mutatis Mutandis: Revisiting the Comparator in Discrimination Testing

### v2

@@ 4. An Illustrative Experiment
$\mathcal{M}\,$$\begin{cases}R&\leftarrow U_{R}\\ G&\leftarrow U_{G}\\ UGPA&\leftarrow b_{U}+\beta_{1}\cdot R+\lambda_{1}\cdot G+U_{1},\\ LSAT&\leftarrow\exp\{b_{L}+\beta_{2}\cdot R+\lambda_{2}\cdot G+U_{2}\}\\ \end{cases}$
$Y$ =$=\begin{cases}1\;\text{if}\;(0.6\cdot UGPA+0.4\cdot LSAT)>\psi\\ 0\;\text{otherwise}\\ \end{cases}$
Figure 1. The causal knowledge with corresponding SCM $\mathcal{M}$ and DAG $\mathcal{G}$ for Law School Admissions. Let $b_{U}$ and $b_{L}$ denote the intercepts; $\beta_{1}$, $\beta_{2}$, $\lambda_{1}$, $\lambda_{2}$ the weights; and $UGPA\sim\mathcal{N}$ and $LSAT\sim\text{Poisson}$ the probability distributions.
To showcase the CP (Definition 2.1) and MM (Definition 3.1) comparators and their impact on discrimination testing, we consider the Law School Admissions example popularized by Kusner et al. (2017) and use on it the discrimination tools of situation testing (ST) (Thanh et al., 2011) and counterfactual situation testing (CST) (Álvarez and Ruggieri, 2023). ST and CST are data mining tools that look for $k$-nearest neighborhoods for complainant-comparator pairs in a dataset. We focus on these two tools because CST extends ST by implementing the MM comparator over the CP comparator. Both tools follow a similar pipeline, even sharing the same distance function for constructing the neighborhoods around the complainant and comparator, with the key difference that ST flips the protected attribute to create its (CP) comparator, whilst CST uses the generated counterfactual of a complainant $i$ to create its (MM) comparator $i^{\prime}$. The code developed for this section is publicly available in a Github repository. [^13] Visit: https://github.com/cc-jalvarez/revisiting-the-comparator.
The dataset contains $n=21790$ applicants, containing the applicants’ gender ($G$), race ($R$), undergraduate grade-point average ($UGPA$), and law school admissions test score ($LSAT$). In terms of the protected attributes, 43.8% of applicants are females and 16.1% are non-whites. Following Álvarez and Ruggieri (2023), we define a decision-maker $b$ that uses only $UGPA$ and $LSAT$ for admitting $Y=1$ or not $Y=0$ an applicant. The decision-maker $b$ test if a weighted sum of $UGPA$ and $LSAT$ is above a cutoff $\psi$, where $\psi$ is the median entry requirement of a top US law school.
We are interested in testing, separately, whether female and non-white applicants are potentially discriminated under this decision-making process. [^14] Intersectionality, meaning considering what occurs with non-white-female applicants, is not considered under most non-discrimination laws (Xenidis, 2020; Crenshaw, 1989). Testing-wise, it means that we can only consider the claims for gender and race separately. To illustrate these tools, consider a female applicant as the complainant $i$. ST finds $k$-females and $k$-males with the closest $UGPA$ and $LSAT$ to $i$ and compare the difference in the proportion of rejections between both groups. The complainant is considered discriminated if such a difference is significantly different from zero at $95\%$ confidence level. Under ST, the ideal comparator $i^{\prime}$ in ST is a male with the same $UGPA$ and $LSAT$ as $i$. CST does the same with the exception that it first generates $i$’s counterfactual using Figure 1 and uses it to find the $k$ males $UGPA$ and $LSAT$. The female group is construct the same way as in ST as the counterfactual generation does not apply to them. Under CST, there is no ideal comparator $i^{\prime}$ to be found but, instead, a counterfactual representation of $i$ in the form of $i^{\prime}$ with $\tilde{\mathbf{X}}$ (6). Because of how each method constructs the male neighborhoods, ST and CST implement, respectively, the CP and MM comparators.
@@
Tables 1 and 2 show the number of discrimination cases identified under ST and CST for, respectively, the protected attributes race (what would have been the outcome had the applicant been white?) and gender (what would have been the outcome had the applicant been male?) for different $k$ neighborhood sizes. [^15] The CST results are without the complainant-comparator pair. This is because ST does not include them, which allows to compare the results for both methods. Results including the pair are similar to those without it. See Álvarez and Ruggieri (2023) for details. Both tables illustrate the impact of choosing between the CP and MM comparators, with CST detecting a higher number of discrimination cases for the same dataset. This difference is driven by ST, under the CP comparator, and CST, under the MM comparator, aiming, respectively for an idealized and a fairness given the difference comparison. Intuitively, under a stricter or more conservative notion of similarity, the ST constructs similar neighborhoods based on the protected attribute. We observe this in Figure 2 that shows seven randomly chosen individuals considered to have been discriminated by both ST and CST on the basis of their gender. We compare the $LSAT$ and $UGPA$ distributions in the “test groups” (i.e., the neighborhoods constructed around the comparator) of ST and CST to the $LSAT$ and $UGPA$’s distribution in the “control group” (i.e., the neighborhood constructed around the complainant), which is the same for both ST and CST. The two groups under ST for $LSAT$ and $UGPA$ are similarly distributed while this is not the case under CST. Figure 2 illustrates $\tilde{\mathbf{X}}$ (6) and its role in discrimination testing. In Figure 3 we carry out the same analysis and find similar results for race.
@@
Under CST, the complainant used for drawing the so-called test groups is purely generated under counterfactual generation. In other words, these profiles determining which male (or white) outcomes are to be compared to the female (or non-white) outcomes are purely hypothetical. Each of these generated profiles represents what would have been of the female (or non-white) individual had it been male (or white) in the admissions process to law school. In generating them and, in turn, adjusting for the impact of gender (or race) on $UGPA$ and $LSAT$, CST creates individual targets for each protected individual that represent where they would be in a fairer (read, free from the impact of the protected attribute) world. CST, thus, tests for equality under these targets. Of course, these results are specific to ST and CST and the potential limitations inherent to both methods, though they still highlight what it would mean to test for discrimination under the MM comparator instead of the CP comparator.
@@
Figure 2. We show the distributions in the form of box-plots for the $LSAT$ and $UGPA$ for seven randomly chosen complainants based on the protected attribute gender. These complainants are found to be discriminated by both ST and CST for $k=100$. The so-called control (ctr) are the neighborhoods of similar female profiles to the complainant, which is the same for both tools. Similarly, the so-called test groups (tst-st for ST; tst-cf for CST) are the neighborhoods of similar male profiles to the comparator: i.e., the CP comparator for ST, and the MM comparator for CST. In both figures, we observe a difference between the two kinds of test groups. As we would expect, it means that the test and control groups are more similar under ST than under CST.
@@
Figure 3. Here, the same analysis follows as in Figure 2 but for the protected attribute race, including similar insights and conclusions on the influence of the CP and MM comparators. In both figures, we observe a difference between the two kinds of test groups.
@@ 5. Overview of Discrimination Testing Tools
In this section, we survey a representative set of tools for discrimination testing, focusing on the comparator. We recommend Romei and Ruggieri (2014) for an extensive and multidisciplinary survey. All these tools amount to gathering prima facie evidence against a decision-maker, meaning they derive the comparator $i^{\prime}$ for the complainant $i$ and compare their outcomes. We view these tools as instances of the counterfactual model of discrimination.
@@ 5.1. Standard Tools
By standard we mean non-algorithmic tools. These tools mainly include natural experiments (e.g., Goldin and Rouse (2000)), field experiments (e.g., Bertrand and Duflo (2017)), and audits (e.g., Fix and Struyk (1993)). These tools share a focus on either gathering or evaluating data that is, by design, able to capture the effect of the protected attribute on the decision outcome. These tools are somewhat outdated for how decisions are taken in modern decision flow.
@@
Let us start with audits, which are a more holistic kind of tool, very human-dependent, and less focused on modeling. Audits consist in examining the decision process in question through human auditors. In practice, it means sending a group of experts to observe during a period of time and study the practices of the decision-maker to corroborate the claims of the complainant (Fix and Struyk, 1993). In the case of the modern decision flow, such a practice only makes sense if the model is used by a human decision-maker or if the focus is to audit the procedure by which the model was trained (Koshiyama et al., 2024), which is increasingly being viewed as a priority over interpretability by regulators (Panigutti et al., 2023).
@@
Natural and field experiments, instead, tend to have more of a focus on discrimination as a modeling problem (i.e., identifying the causal effect of the protected attribute) and, thus, prioritize designing and gathering data suitable for this purpose. These methods are a product of their time. In particular, these methods are a product of the so-called “empirical revolution” experienced by the social sciences in which a mixture of data availability, better computers, and an intuitive causal framework (the Potential Outcomes Framework) made it possible to test causal claims via experiments (Angrist and Pischke, 2008).
@@
The overall idea behind natural experiments is simple. A policy introduction (or external shock) occurs creating a before-and-after scenario that, in principle, allows us to test for the effect of the protected attribute and find ourselves, naturally, in an experimental setting. Goldin and Rouse (2000)’s Orchestrating Impartiality: The Impact of“Blind” Auditions on Female Musicians is a clear example. US orchestras before the 1970s used to audition musicians without screens, meaning musicians would perform in full sight to the evaluating committee. The US orchestra system back then were male dominated and, some, openly sexist, which motivated the suspicion that there was discrimination against female musicians. Orchestras then started to conduct blind auditions to ensure an impartial process. Some orchestras even introduced carpets so that female musicians would not be identified by the sound of their heels when walking up the stage. Female representation in orchestras increased after the introduction of the screen. Goldin and Rouse (2000), though, using data from the auditions, set out to study whether this amounted to evidence for discrimination within the US orchestra system. There was indeed a before-and-after mechanism through the introduction of the screen that coincided with an increase in female musicians in orchestras, but to claim a causal effect we needed to control for other factors. For instance, the introduction of the screen also coincided with an overall increase of females in the US work force. Goldin and Rouse (2000) used a regression model to control for this and other factors, finding inconclusive evidence for discrimination against female musicians. Goldin and Rouse (2000) is a seminal paper, among other things, because it shows how meticulous the model choice has to be to claim discrimination evidence. It illustrates, implicitly, the importance of deriving a comparator good-enough to isolate the effect of the protected attribute.
@@
Field experiments, instead, focus on designing an experiment to create the proper data for testing the discrimination claim. The most famous example is Bertrand and Mullainathan (2004)’s Are Emily and Greg More Employable Than Lakisha and Jamal? A Field Experiment on Labor Market Discrimination. Also see Bertrand and Duflo (2017) for a survey on field experiments for discrimination. Bertrand and Mullainathan (2004) proposed the following experiment: let us create pairs of fictitious CVs tailored to job positions that are the same except for the candidate’s name, where the name is intended to reflect both gender and race within the US context. They would create, e.g., the same CV for Mike and Jamal (respectively, a common white and black name according to census data at the time); send both CVs to the same job openings via mail (this was done in the early 2000s); and compare which CV received a callback. The CV with the black-sounding name was the complainant and the CV with the white-sounding name its comparator. All the CVs were drafted by the researchers. In the end, they found evidence for discrimination. This sort of field experiment is known as a correspondence study. See Rooth (2021) for a survey on correspondence studies. Another kind of field experiment is situation testing (Rorive, 2009) in which, rather than sending CVs or other forms of correspondence, we send a group of actors that share key physical traits and backstories as the complainant and its comparator.
@@
Standard tools are at the center of the criticism against the counterfactual model of discrimination. This fact is not surprising, as these tools have been around longer, even inspiring the recent algorithmic tools. Heckman (1998), e.g., is critical of the effectiveness of these methods in controlling for similarity and other factors and argues that we have been too confident in claiming to be able detect discrimination. More recently, Kohler-Hausmann (2019) goes after these methods (she is particularly critical of Bertrand and Mullainathan (2004)) due to their experimental designs. These standard tools rely on the CP comparator. As previously discussed, under direct discrimination, these tools are still useful and have a lot to offer to algorithmic tools in terms of their focus on inferential statistics. If, however, we are interested in proving indirect discrimination, then these experiments have to make a strong argument for using the CP comparator. In particular, this argument should be centered on the kind of signal (as in relevant information) being used by the decision-maker and whether that signal has any link to the protected attribute. Consider the case of the orchestra auditions. There, the decision-maker cares only for the playing quality of the musician, which should not be related to gender. [^16] Though there is a reasonable case to be made here that instruments tend to favor male musicians (Criado-Perez, 2019). Hence, the signal can be recovered once the screen is imposed as the sounds can go through the screen but all other factors, like the musician’s appearance, are blocked. Consider now the case of the CVs. There, even if we impose a “screen” by removing the names, there is still the possibility that the decision-maker uses signals for candidate potential that are also linked to race or gender, such as extracurricular activities or the name of the high-school. These are valid signals, or at least signals that could be well argued by the decision-maker as necessary for measuring candidate potential. In such a setting, the MM comparator might be better suited.

### v3

@@ 4. An Illustrative Experiment
To showcase the CP (Definition 2.1) and MM (Definition 3.1) comparators and their impact on discrimination testing, we consider the Law School Admissions example popularized by Kusner et al. (2017) and use on it the discrimination tools of situation testing (ST) (Thanh et al., 2011) and counterfactual situation testing (CST) (Álvarez and Ruggieri, 2023, 2025). ST and CST are both data mining tools that look for $k$-nearest neighborhoods for complainant-comparator pairs in a dataset. We focus on these tools because CST extends ST by implementing the MM comparator over the CP comparator. Both tools follow a similar pipeline, even sharing the same distance function for constructing the neighborhoods around the complainant and comparator, with the key difference that ST flips the protected attribute to create its (CP) comparator while CST uses the generated counterfactual of a complainant to create its (MM) comparator. CST does so given the SCM $\mathcal{M}$ with DAG $\mathcal{G}$ shown in Figure 1. The code for this section is publicly available in a GitHub repository. [^14] https://github.com/cc-jalvarez/revisiting-the-comparator.
The dataset contains $n=21790$ applicants, containing the applicants’ gender ($G$), race ($R$), undergraduate grade-point average ($UGPA$), and law school admissions test score ($LSAT$). In terms of the protected attributes, 43.8% of applicants are females and 16.1% are non-whites. Following Álvarez and Ruggieri (2023, 2025), we define a decision-maker $b(\cdot)$ that uses only $UGPA$ and $LSAT$ for admitting $Y=1$ or not $Y=0$ an applicant. The decision-maker $b(\cdot)$ tests if an applicant’s weighted sum of $UGPA$ and $LSAT$ is above the cutoff $\psi$, where $\psi$ is the median entry requirement of a top US law school. [^15] The cutoff is the weighted sum of 60% in $UGPA$ (3.93 over 4.00), and 40% $LSAT$ (46.1 over 48), giving total of 20.8. The maximum possible score given $b(\cdot)$ is 22. It is based on Yale University Law School: https://www.ilrg.com/rankings/law/index/1/asc/Accept. We are interested in testing, separately, whether female and non-white applicants are potentially discriminated under this decision-making process. [^16] Intersectionality, meaning what occurs with non-white-female applicants, is not considered under most non-discrimination laws (Xenidis, 2020; Crenshaw, 1989; Álvarez and Ruggieri, 2025). Testing-wise, it means that we only consider the claims for gender and race separately.
CST proceeds as follows. Given a tabular dataset of decisions, with each row denoting an individual, it detects whether each protected individual was or not (potentially) discriminated by the decision-maker. Each protected individual represents a complainant $i$, and for each $i$ it generates its comparator $i^{\prime}$ by running a counterfactual query on the protected attribute of interest (recall, Section 1.1) based on a SCM. Using Gower’s distance (Gower, 1971), CST mines for $k$-similar protected individuals to $i$ and $k$-similar non-protected individuals to $i^{\prime}$. The former represents the control group while the latter represents the test group. Recall that $i^{\prime}$ does not appear in the input dataset and answers to the question what would have been of $i$ had it been a non-protected individual during the decision process? The control group is centered on $i$ while the test group is centered on $i^{\prime}$. CST then compares the difference in rejection proportions—meaning how many individuals out of $k$ in a group received a negative decision—between the control and test groups. In its strictest form, CST considers as evidence for discrimination if the difference in proportions is strictly greater than zero, meaning the proportion of negative outcomes is greater in the control group compared to the test group. Because it evaluates a difference in proportions, CST is able to equip each individual claim with confidence intervals. We do not use this aspect of CST here. We include the relevant details of CST in Appendix B.
ST follows similarly, with the key exception that it does not generate the complainant $i^{\prime}$. Instead, it builds both control and test groups around $i$ using Gower’s distance (Gower, 1971). It means that $i^{\prime}$, in practice, is the most similar non-protected individual to $i$ according to the distance function. Both control and test groups are centered around $i$. For our purposes in this section, the rest of the procedure is similar to CST. Importantly, ST excludes the complainant-comparator pair when looking at the difference in proportion of negative outcomes. CST, instead, has the option to include or exclude the “search centers”.
{tikzpicture}
@@
[…]
$\mathcal{M}\,$ $\begin{cases}R&\leftarrow U_{R}\\ G&\leftarrow U_{G}\\ UGPA&\leftarrow b_{U}+\beta_{1}\cdot R+\lambda_{1}\cdot G+U_{1}\\ LSAT&\leftarrow b_{L}+\beta_{2}\cdot R+\lambda_{2}\cdot G+U_{2}\end{cases}$
[…]
Figure 1. The auxiliary causal knowledge with corresponding SCM $\mathcal{M}$ and DAG $\mathcal{G}$ for Law School Admissions (Level 3) (Kusner et al., 2017). Let $b_{U}$ and $b_{L}$ denote the intercepts and $\beta_{1}$, $\beta_{2}$, $\lambda_{1}$, $\lambda_{2}$ the regression weights. These parameters are estimated from the data using ordinary linear regression. We obtain, respectively, the estimates $\hat{b_{U}}=3.21$, $\hat{\beta_{1}}=-0.22$, $\hat{\lambda_{1}}=0.13$ and $\hat{b_{L}}=37.8$, $\hat{\beta_{2}}=-4.64$, $\hat{\lambda_{2}}=-0.61$. :
To illustrate these two tools, consider a female applicant as the complainant $i$. ST finds $k$-females and $k$-males with the closest $UGPA$ and $LSAT$ to $i$ and compares the difference in the proportion of rejections between both groups. The complainant is considered discriminated if such a difference is significantly different from zero. The ideal comparator $i^{\prime}$ in ST is a male with the same $UGPA$ and $LSAT$ as $i$. CST does the same with the exception that it first generates $i$’s counterfactual via Figure 1 and uses it to find $k$-males based on the closest $UGPA$ and $LSAT$. The female group is constructed in the same way as in ST as the counterfactual generation does not apply to them. This is because the counterfactual query of interest (what would have happened to the female applicant had she been male?) is concerned with the group of male applicants. There is no ideal comparator $i^{\prime}$ to be found under CST but, instead, a counterfactual representation of $i$ in the form of $i^{\prime}$ with $\tilde{\mathbf{X}}$ (6).
@@
Because of how each method constructs the male neighborhoods, ST and CST implement, respectively, the CP and MM comparators. Tables 1 and 2 show the number of prima facie discrimination cases identified under ST and CST for, respectively, the protected attributes race (with the goal of answering what would have been the outcome had the applicant been white?) and gender (and what would have been the outcome had the applicant been male?) for different $k$ neighborhood sizes. CST generates a counterfactual distribution for each query. The CST results are without the complainant-comparator pair. This is because ST does not include them. We run CST without these “search centers” to have comparable results between both methods.
@@
The tables illustrate the impact of choosing between the CP and MM comparators, with CST detecting a higher number of prima facie discrimination cases for the same dataset. This difference is driven by ST, under the CP comparator, and CST, under the MM comparator, aiming, respectively, for an idealized and a fairness given the difference comparison. Intuitively, under a stricter or more conservative notion of similarity, the ST constructs similar neighborhoods based on the protected attribute. For instance, we observe this difference in Figure 2 that shows seven randomly chosen individuals considered to have been discriminated by both ST and CST on the basis of their gender. We compare the $LSAT$ and $UGPA$ distributions in the “test groups” (i.e., the neighborhoods constructed around the comparator) of ST and CST to the $LSAT$ and $UGPA$’s distribution in the “control group” (i.e., the neighborhood constructed around the complainant), which is the same for both ST and CST. The two groups under ST for $LSAT$ and $UGPA$ are similarly distributed while this is not the case under CST. Figure 2 illustrates $\tilde{\mathbf{X}}$ (6) and its role in discrimination testing. In Figure 3 we carry out the same analysis and find similar results for race.
The results are starker for race than for gender. This point is observed by the difference between ST and CST for each protected attribute across $k$ size. For instance, at $k=50$, CST almost triples ST in number of detected cases for race but shows just a marginal increase for gender. What is clear from these results is that the effect of race and gender on $LSAT$ and $UGPA$, within at least the context of law school admissions, is not symmetrical—as also confirmed by the parameter estimates reported in the caption of Figure 1. As argued in other works (Fryer Jr and Torelli, 2010; Austen-Smith and Fryer, 2005), each protected attribute conditions differently its members as each protected attribute came to be, at least under a social construct perspective (Mallon, 2007), with different purposes. Based on these results and our conceptions of the CP and MM comparators, the similar results between ST and CST for gender would indicate a less pronounced impact from gender on $LSAT$ and $UGPA$. In other words, not much changes when generating the counterfactual profiles. The opposite occurs with race. These results seem in line with work on how race can be more detrimental than gender in US college admissions (Baker et al., 2018; Conger and Long, 2013). Further, these results point at CST’s capacity to capture less pronounced forms of discrimination. Be that as it may, both tables do show evidence of potential discrimination by the decision-maker $b(\cdot)$, with CST detecting more cases. As discussed in Section 2.2, detecting discrimination is one aspect of establishing discrimination. Indeed, CST paints a worst setting than ST for the admissions process in question, but, regardless, the results of ST or CST would then have to be argued among stakeholders (such as the university itself). It varies on a case-by-case basis and the university can always provide, in principle, an objective justification.
@@
The complainant used by CST for drawing the so-called test groups is generated under counterfactual generation using the abduction, action, and prediction steps (Pearl et al., 2016). These profiles determining which male (or white) outcomes are to be compared to the female (or non-white) outcomes are purely hypothetical. Each of these generated profiles represents what would have been of the female (or non-white) individual had it been male (or white) in the admissions process to law school. In generating them and, in turn, adjusting for the impact of gender (or race) on $UGPA$ and $LSAT$, CST creates individual targets for each protected individual that represent where they would have been in a fairer (read, free from the impact of the protected attribute) world. CST tests for equality under these targets.
These results are specific to ST and CST and the potential limitations inherent to both methods, though they still highlight what it would mean to test for discrimination under the MM comparator instead of the CP comparator. Notably, the notion of generating the comparator is non-trivial. A method like ST is more conservative in its treatment of similarity between protected and non-protected individuals but, at the same time, it is also more intuitive. The CP comparator is the standard because it offers a simple view on similarity. A method like CST, instead, enters into a more subjective realm of what similarity should be. Although it relies on a modeling counterfactual approach in which the generated counterfactual distribution is unique to the given SCM and the specific intervention, said distribution that support the instances used as comparators is one of other plausible counterfactual representations of how the world should be (Woodward, 2005). Some representations are more plausible than others, but still it is not as though CST builds on a shared groundtruth of what the world should be. The MM comparator, in that regard, poses a more challenging and, even, uncomfortable challenge of picturing in practice the sort of equality we want to aim for (Wachter et al., 2020).
Further, in the case of CST, it is worth noting that the derivation of the SCM in Figure 1 is a difficult task. For our purposes, which is to showcase the CP versus MM comparison, this limitation is taken for granted. However, in practice, it requires either agreeing on the SCM for a given context, or discovering it based on data and knowledge using causal discovery techniques (Vowels et al., 2023; Nogueira et al., 2022). Álvarez and Ruggieri (2025) provide a discussion on this limitation, which, in short, can be summarized as a pragmatic view on causal modeling (Loftus, 2024): the SCM help to explicitly state our assumptions about the world. That is what makes them useful in setting in which disagreement among stakeholders is likely, such as discrimination testing. Additionally, CST depends on a SCM that describes the dataset and the complexity of the dataset used will, thus, condition the SCM. Similarly, CST depends on Gower’s distance function. Here, we did not explore these limitations. These limitations are independent of the MM comparator as a concept, though should be taken into account in case we wish to implement it using SCM.
@@
Figure 2. We show the distributions in the form of box-plots for the $LSAT$ and $UGPA$ for seven randomly chosen complainants based on the protected attribute gender. These complainants are found to be discriminated by both ST and CST for $k=100$. The so-called control (ctr) are the neighborhoods of similar female profiles to the complainant, which is the same for both tools. Similarly, the so-called test groups (tst-st for ST; tst-cf for CST) are the neighborhoods of similar male profiles to the comparator: i.e., the CP comparator for ST, and the MM comparator for CST. In both figures we observe a difference between the two kinds of test groups. As we would expect, it means that the test and control groups are more similar under ST than under CST. Both figures show the impact of the CP and MM comparators. :
@@
Figure 3. The same analysis follows as in Figure 2 but based on the protected attribute race, including similar insights and conclusions on the influence of the CP and MM comparators. In both figures, we observe a difference between the two kinds of test groups, highlighting again the impact of the CP and MM comparators. :
@@ 5. Overview of Discrimination Testing Tools
In this section, we survey a representative set of tools for discrimination testing focusing on the comparator. Different from Section 1.2, the goal is not to position this work but rather to provide an overview of the methods, which span multiple fields, while using our novel formalization of the ceteris paribus and mutatis mutandis comparators. We recommend Romei and Ruggieri (2014) for an extensive and multidisciplinary survey. All these tools amount to gathering prima facie evidence against a decision-maker, meaning they derive the comparator $i^{\prime}$ for the complainant $i$ and compare their outcomes. We view these tools as instances of the counterfactual model of discrimination.
@@ 5.1. Standard Tools
Here, by standard we mean non-algorithmic tools. These tools mainly include natural experiments (e.g., Goldin and Rouse (2000)), field experiments (e.g., Bertrand and Duflo (2017)), and audits (e.g., Fix and Struyk (1993)). These tools share a focus on either gathering or evaluating data that is, by design, able to capture the effect of the protected attribute on the decision outcome. Such data that is characterized by a before-and-after or a with-and-without mechanism around the protected attribute that ensures the identifiability of the protected attribute’s effect. These tools are somewhat outdated for how decisions are taken in modern decision flows, in which ML may enable and even automate key steps.
@@
Let us start with audits, which are a more holistic kind of tool, very human-dependent, and less focused on modeling. Audits consist in examining the decision process in question through human auditors. In practice, it means sending a group of experts to observe during a period of time the decision-maker and study its practices to corroborate the claims of the complainant (Fix and Struyk, 1993). In the case of the modern decision flow, such a practice only makes sense if the model is used by a human decision-maker or if the focus is to audit the procedure by which the model was trained (Koshiyama et al., 2024), which is increasingly being viewed as a priority over interpretability by regulators (Panigutti et al., 2023). The focus is on understanding the overall process and identifying procedural concerns. The comparator is not an explicit objective.
@@
Natural and field experiments, instead, have more of a focus on discrimination as a modeling problem (i.e., identifying the causal effect of the protected attribute) and, thus, prioritize the designing and gathering of data suitable for this purpose. Notably, these methods are a product of their time. In particular, these methods are a product of the so-called “empirical revolution” experienced by the social sciences in which a mixture of data availability, faster and cheaper computers, and an intuitive causal framework—the Potential Outcomes Framework—made it possible to test causal claims via experiments (Angrist and Pischke, 2008).
@@
The overall idea behind natural experiments is simple. A policy introduction or external shock occurs, creating a before-and-after scenario that, in principle, allows us to test for the effect of the protected attribute. We, thus, find ourselves naturally in an experimental setting. Goldin and Rouse (2000)’s Orchestrating Impartiality: The Impact of “Blind” Auditions on Female Musicians is a clear example. US orchestras before the 1970s used to audition musicians without screens, meaning musicians would perform in full sight to the evaluating committee. The US orchestra system back then was male dominated and openly sexist, which motivated the suspicion that there was discrimination against female musicians. Orchestras then started to conduct blind auditions to ensure an impartial process. Some orchestras even introduced carpets so that female musicians would not be identified by the sound of their heels when walking up the stage. Female representation in US orchestras increased after the introduction of the screen. Goldin and Rouse (2000), though, using data from the auditions, set out to study whether this amounted as evidence for gender discrimination within the US orchestra system. There was indeed a before-and-after mechanism through the introduction of the screen that coincided with an increase in female musicians in orchestras, but to claim a causal effect it was necessary to control for other factors. For instance, the introduction of the screen also coincided with an overall increase of female participation in the US work force. Goldin and Rouse (2000) used a regression model to control for this and other factors, finding inconclusive evidence for discrimination against female musicians. Goldin and Rouse (2000) is a seminal paper, among other things, because it shows how meticulous the model choice has to be to claim discrimination evidence. The paper illustrates the importance of deriving a comparator good-enough to isolate the effect of the protected attribute.
@@
Field experiments, instead, focus on designing an experiment that creates the data for testing the discrimination claim. The most famous example is Bertrand and Mullainathan (2004)’s Are Emily and Greg More Employable Than Lakisha and Jamal? A Field Experiment on Labor Market Discrimination. Also see Bertrand and Duflo (2017) for a survey on field experiments for discrimination. Bertrand and Mullainathan (2004) proposed the following experiment: let us create pairs of fictitious CVs tailored to job positions that are the same except for the candidate’s name, where the name is intended to reflect both gender and race within the US context. They would create, for example, the same CV for Mike and Jamal (respectively, a common white and black name according to census data at the time); send both CVs to the same job openings via mail (this was done in the early 2000s); and compare which CV received a callback. The CV with the black-sounding name was the complainant and the CV with the white-sounding name its comparator. All the CVs were drafted by the researchers. In the end, they found evidence for discrimination. This sort of field experiment is known as a correspondence study. See Rooth (2021) for a survey on correspondence studies. Another kind of field experiment is situation testing (Rorive, 2009) in which, rather than sending CVs or other forms of correspondence, we send a group of actors that share key physical traits and backstories as the complainant and its comparator.
Standard tools are at the center of the criticism against the counterfactual model of discrimination. This fact is not surprising, as these tools have been around longer, even inspiring the recent algorithmic tools in the next sections. They represent the starting point for more novel discrimination testing methods, such as situation testing (Thanh et al., 2011) and counterfactual situation testing (Álvarez and Ruggieri, 2025) both proposing an algorithmic version of the human-based situation testing (Rorive, 2009). Heckman (1998), for example, is critical of the effectiveness of these methods in controlling for similarity and other factors and argues that we have been too confident in claiming to be able detect discrimination. More recently, as we discussed in earlier sections, Kohler-Hausmann (2019) attacks these methods (she is particularly critical of Bertrand and Mullainathan (2004)) due to their experimental designs and causal treatment of the protected attribute.
@@
These standard tools rely on the CP comparator. As discussed previously, under direct discrimination, these tools are still useful and still have a lot to offer to algorithmic tools in terms of their focus on inferential statistics. If, however, we are interested in proving indirect discrimination, then these tools and their experiments need to make a stronger argument for using the CP comparator. In particular, this argument should be centered on the kind of signal, as in relevant information, being used by the decision-maker and whether that signal has any link to the protected attribute in question. Consider the case of the orchestra auditions (Goldin and Rouse, 2000). There the decision-maker cares only for the playing quality of the musician, which should not be related to gender. [^17] Though there is a reasonable case to be made that instruments tend to favor male musicians (Criado-Perez, 2019). Hence, the signal can be recovered once the screen is imposed as the sounds can go through the screen but all other factors, like the musician’s appearance, are blocked. In such a setting, the CP comparator might already be well suited. Consider now the case of the CVs (Bertrand and Mullainathan, 2004). There, even if we impose a “screen” by removing the names, there is still the possibility that the decision-maker uses signals for candidate potential that are linked to race or gender, such as extracurricular activities or the name of the high-school. These are valid signals, or at least signals that could be well argued by the decision-maker as necessary for measuring candidate potential. In such a setting, the MM comparator might be better suited.